The band members sold tickets to their concert performance. Some were $5 tickets, and some were $6 tickets. There were 16 more than twice as many $6 tickets sold as $5 tickets. The total sales were $1643. How many tickets of each price were sold?
That is the problem, but I just need the right formula to go with it. Thank you.

Answer by mananth(16946)

x= Number of $5 tickets
y= Number of $6 tickets
Total sold value
5 x + 6 y = 1643 .............1
condition
y=2x+16
-2 x + 1 y = 16 .............2
Eliminate y
multiply (1)by -1
Multiply (2) by 6
-5 x -6 y = -1643
-12 x + 6 y = 96
Add the two equations
-17 x = -1547
/ -17
x = 91
plug value of x in (1)
5 x + 6 y = 1643
455 + 6 y = 1643
6 y = 1643 -455
6 y = 1188
y = 198
x= 91 Number Of $5 tickets
y= 198 Number Of $6 tickets
m.ananth@hotmail.ca